OpenStudy (anonymous):

PLEASE HELP! Determine if the triangle with the following vertices is a right triangle. A(-3, 5), B(1, 4), and C(-1, 0). tell whether or not the triangle is a right triangle. Explain how you arrived at your answer.

OpenStudy (anonymous):

What are you stuck on?

OpenStudy (campbell_st):

start by plotting the points... then decide which sides may intersect at right angles... then find the slopes of the 2 sides to check. remember the product of the slopes needs to be -1 for 2 lines to be perpendicular

OpenStudy (anonymous):

I dont have any graph paper to plot it on

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... just sketch them  so it looks like AC and BC might be perpendicular.. so find the slopes and check

OpenStudy (wolf1728):

Reloading that posting.

OpenStudy (wolf1728):

Slope AB = -1/4 Slope BC = 2 So BC and AB are not perpindicular

OpenStudy (anonymous):

So it isnt a right triangle correct?

OpenStudy (wolf1728):

Or to analyze it further Line AB² = 4² +1² Line AB² = 17 Line AB = 4.1231056256 Line BC² = 4² + 2² Line BC² = 20 Line BC = 4.472135955 Line AC² = 5² + 2² Line AC² = 29 Line AC = 5.3851648071 If AB and BC are perpendicular, then AB² + BC² must equal AC² but they don't.

OpenStudy (anonymous):

Thank you!